In the mid-1990s, time magazine reported that 27% of the US Congress supported a tax cut as a means of stimulating the economy and increasing Tax revenues. Suppose at that time five members of congress were randomly selected for an interview and asked whether they supported the tax cut to stimulate the economy.

Find the probability that:

2.1 At least three of the five members were in favour of a tax cut. (6)

2.2 Less than five members were in favour of a tax cut. (6)

2.3 Not more than 3 members were in favour of the tax. (6)

2.4 Find the mean of the above distribution (6)

2.5 Find the variance and standard deviation of the above distribution (6)


1
Expert's answer
2022-04-19T03:29:14-0400

Let "p=0.27,~q=1-p=0.73,~n=5"

2.1) According to the binomial distribution:

"P(x\ge3)=P(x=3)+P(x=4)+P(x=5)=\\
=\sum_{k=3}^{5}C_n^kp^kq^{n-k}=\frac{5\cdot4\cdot3}{1\cdot2\cdot3}\cdot0.27^3\cdot0.73^2+\\
+\frac{5\cdot4\cdot3\cdot2}{1\cdot2\cdot3\cdot4}\cdot0.27^4\cdot0.73+0.27^5=\\
=0.125723194\approx12.57\%"


"2.2)~P(x<5)=1-P(x=5)=\\
=1-C_5^5p^5q^0=1-0.27^5=\\
=0.998565109\approx99.86\%"


"2.3)~P(x\le3)=1-P(x=4)-P(x=5)=\\
=1-5\cdot0.27^4\cdot0.73-0.27^5=\\
=0.979167513\approx97.92\%"


"2.4)~\mu=\sum_{k=0}^nkC_n^kp^kq^{n-k}=\\
=\sum_{k=1}^nk\frac{n!}{k!(n-k)!}p^kq^{n-k}=\\
=\sum_{k=1}^n\frac{n(n-1)!}{(k-1)!((n-1)-(k-1)!}pp^{k-1}q^{(n-1)-(k-1)}=\\
=np\sum_{k=0}^{n-1}C_{n-1}^kp^kq^{n-1-k}=\\
=np(p+q)^{n-1}=np=5\cdot0.27=1.35"


"2.5)~\sigma^2=\frac{1}{n-1}\sum_{k=0}^nC_n^kp^kq^{n-k}(k-\mu)^2=\\
=\frac{1}{4}(0.73^5\cdot1.35^2+5\cdot0.27\cdot0.73^4\cdot0.35^2+\\
+\frac{5\cdot4}{1\cdot2}\cdot0.27^2\cdot0.73^3\cdot0.65^2+\\
+\frac{5\cdot4\cdot3}{1\cdot2\cdot3}\cdot0.27^3\cdot0.73^2\cdot1.65^2+\\
+5\cdot0.27^4\cdot0.73\cdot2.65^2+0.27^5\cdot3.65^2)=\\
=0.246375\approx0.246\\
\sigma=\sqrt{\sigma^2}=0.496361763\approx0.496"